OpenStudy (anonymous):

what multiplys to 225 but adds to 30

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf 15\times 15 }\) \(\Huge\color{blue}{ \sf 15 + 15 }\)

OpenStudy (mathmale):

I think you mean: "Which two numbers have a product of 225 and a sum of 30?" This is definitely an algebra problem. Please choose a literal (a letter, such as x or y) to represent each of the two unknown numbers. Then write equations in terms of those literals for the product and sum. Solve the 2 equations simultaneously.

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf | 15 |\times| 15 | }\) \(\Huge\color{blue}{ \sf | 15 |+ | 15 | }\) I factored out 225 \(\Huge\color{blue}{ \sf 225=45\times 5=25\times 9== 15\times 15 }\)

OpenStudy (solomonzelman):

I am joking about the absolute value it's the same thing :)

OpenStudy (solomonzelman):

so choosing out of those 3 wouldn't be hard

OpenStudy (mathstudent55):

@mathmale I never heard the term "literal." Is that the same as a variable?

OpenStudy (mathmale):

After double-checking on the meaning of the word "literal," I've decided to "recall" that word and substitute "letter" or "letter of the alphabet." :) In either case, the idea is that we pick a symbol, almost always a letter of the alphabet, to represent each of the unknowns in an algebra problem. Appreciate your asking for clarification!

OpenStudy (mathmale):

I think you mean: "Which two numbers have a product of 225 and a sum of 30?" This is definitely an algebra problem. Please choose a literal (a letter, such as x or y) to represent each of the two unknown numbers. Then write equations in terms of those literals for the product and sum. Solve the 2 equations simultaneously. Let the first number be x and the second number be y. The product of the numbers is 225, so we'd write x*y=225. The sum of the numbers is 30, so we'd write ... what? Then we'd solve these two equations simultaneously to find the values of x and y, which would be the same thing as solving for the "two numbers." Good luck!